We have put in a new EPICS input into the SUS library part, just before the OL_PIT and OL_YAW filter banks, so that the IN1 point is calibrated to microradians. I recompiled all SUS-related models. The OPTLEV_SERVO screen has been changed, so that you can see the calibration, and enter a value. The gains have been reduced by a factor reciprocal to the calibration, so the loop gain is the same.
ETMs, SRM and MCs all have "calibration" numbers of 1, so the numbers aren't really calibrated, they're just the same as always.
It looks like the PRM and the BS are moving significantly (factor of ~30) more than the ITMs at a few Hz! (Y-axis of plot is urad/rtHz)
EDIT JCD: We need to fix up the MEDM QPD indicators, and the OpLev red lights on the watchdog screen, so they match the new numbers. Also, Rana turned on the output limiters to 2000 for all oplev servos.